### v1

# Bounds on the photon mass via the Shapiro effect in the solar system

Abstract
We study the effects of a finite mass for the photon on its propagation in a weak gravitational field. In particular, we analyse the gravitational time delay, also known as the Shapiro effect. We work in isotropic coordinates in the weak-field limit and find that the mass-dependent corrections enhance the gravitational time delay. Doppler-tracking data from the Cassini mission allow us to set upper bounds on the photon mass at the level of $m_{\gamma}\lesssim 3\times 10^{-7}\,{\rm eV/c^{2}}$. We also discuss next-generation solar-system tests of general relativity that could improve this upper limit, potentially reaching $m_{\gamma}\lesssim 8\times 10^{-8}\,{\rm eV/c^{2}}$. Though not competitive with the currently best limits, our projected bounds are at the ballpark of earlier ones based on the gravitational deflection angle.
I Introduction
The photon is the paradigm for a massless particle. In fact, the invariance of electromagnetism under local gauge transformations requires $m_{\gamma}=0$, though the photon may indeed acquire an effective mass in certain circumstances Skobelev ; Cover ; supercond . The fundamental masslessness of the photon is a theoretically pleasing hypothesis, usually taken for granted, but it must be thoroughly tested. The most stringent upper bound currently accepted is $m_{\gamma}\leq 10^{-18}\,{\rm eV/c^{2}}\approx 10^{-51}$ g extracted from data from the Voyager missions Ryutov ; for reviews, see the latest edition of the Particle Data Group PDG , as well as refs. Tu ; Nieto ; Okun ; Goldhaber ; Goldhaber2 .
Measuring the photon mass is only possible if the dimensions of the system are very large or the measurement is very precise Goldhaber ; Goldhaber2 ; Nieto , since the photon mass only appears squared and effects are $\mathcal{O}(m_{\gamma}^{2}d^{2})$, where $d$ is the typical dimension of the system. This severely limits the range of feasible tests, since experiments involving atomic physics, which reach astounding precision, are bound to atomic dimensions, thus leading to relatively weak upper limits Accioly2 . At the intermediate planetary scale the limits profit from larger length scales, but still suffer from relatively low-precision, or noisy data Kroll ; Malta . Finally, at the other extreme, solar-system tests benefit from astronomical distances and place the currently strongest upper bounds Ryutov ; Retino .
Here we study the effect of the photon mass on the gravitational time delay, first suggested by Shapiro in 1964 as a “fourth test of general relativity” Shapiro1 . His idea was to send radio signals from radio telescopes on Earth towards Mercury or Venus when these are close to superior conjunction, i.e., passing almost behind the Sun. According to Shapiro’s calculations, the round-trip time of the signal would be longer than expected in the absence of the Sun: for grazing incidence at the solar limb, the extra travel time due to gravity can reach $\sim 200\,{\rm\mu s}$ in the inner solar system.
The measurements were conducted by Shapiro et al. Shapiro2 ; Shapiro3 using the Arecibo and Haystack telescopes with dedicated radio systems, which allowed the detection of the extremely faint echoes ($\approx 10^{27}$ times weaker than the emitted signals). Curvature effects in general relativity are characterized by Eddington’s $\gamma$ parameter, which is unity in general relativity Poisson ; Will_book , and the data analysis yielded $\gamma=1.03\pm 0.04$, thus confirming the predictions of general relativity to within a few percent. Other similar analyses were undertaken using ranging data from the Mariner landers Mariner1 and the Viking spacecraft Viking ; Viking2 ; Viking3 , both sent to Mars, determining $\gamma$ to within $3\%$ and $0.2\%$ accuracy, respectively. The most precise determination of $\gamma$ so far stems from Doppler-tracking data from the Cassini spacecraft on its way to Saturn, which yielded $|\gamma-1|=2.3\times 10^{-5}$ Cassini_1 ; Cassini_2 . These results will be explored in Sec. III.3 to set upper bounds on the photon mass.
Quite generally, the search for deviations of general relativity, or new-physics effects in general, can profit greatly from experiments in space Book_space . Tests of the universality of free fall, for example, may be performed over longer time periods, thus enabling higher precision levels to be reached. Similarly, interferometric experiments can benefit from much longer interaction times, as well as longer distances, both helping to enhance possible signals. The absence of seismic vibrations and atmospheric disturbances, such as optical distortions and drag (which imparts non-gravitational accelerations to sensors) are further advantages of space-based experiments. Also for time-delay measurements the space environment is ideal, since precise time keeping is crucial when minute height and time inaccuracies may impact the results. In space these variables can be particularly well controlled, thus allowing for very precise measurements.
So far all measurements of the Shapiro effect were performed using ground-based antennas sending and receiving signals to spacecraft at interplanetary distances. Now, however, the technologies necessary to measure time delays to picosecond accuracy in space are mature (e.g., highly stable laser sources, ultra-precise atomic clocks, drag-free control, etc Book_space ). In this context, we also analyse a few promising space missions which include time-delay measurements as their scientific targets, aiming at drastically improving the current limits on $|\gamma-1|$. The projected uncertainties in $\gamma$ will be used to derive sensitivities to the photon mass, cf. Sec. III.4.
This paper is organized as follows: in Sec. II we set up the theoretical basis of our analysis and find the gravitational time delay in the case of massive photons. Next, in Sec. III we apply our result to current and future limits on $|\gamma-1|$ to extract upper bounds on the photon mass. Finally, in Sec. IV we present our conclusions. We use SI units throughout.
II Gravitational effects
We are interested in the behavior of massive Abelian spin-1 bosons propagating in a weak gravitational field. The dynamics of the massive photon is governed by the de Broglie-Proca action coupled to gravity dB1 ; dB2 ; dB3 ; Proca1 ; Proca2
$S=\int d^{4}x\sqrt{-g}\left[-\frac{1}{4\mu_{0}}F^{\mu\nu}F_{\mu\nu}+\frac{\mu_{\gamma}^{2}}{2\mu_{0}}A_{\mu}A^{\mu}\right]\,,$ (1)
where $\mu_{\gamma}=m_{\gamma}c/\hbar$ is the reciprocal reduced Compton wavelength and $\mu_{0}$ is the vacuum permeability. The anti-symmetric field-strength tensor is $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ and $g$ is the determinant of the metric tensor.
Before starting we must decide on the coordinates to describe the spacetime geometry of a point mass. While standard Schwarzschild coordinates are common, here we employ Sun-centered isotropic spherical coordinates with which the interval squared becomes
$ds^{2}=\left(\frac{1-\frac{GM}{2c^{2}r}}{1+\frac{GM}{2c^{2}r}}\right)^{2}c^{2}dt^{2}-\left(1+\frac{GM}{2c^{2}r}\right)^{4}\left(dr^{2}+r^{2}d\Omega^{2}\right)$ (2)
[…]
Isotropic coordinates are typically used in astrometric calculations and are particularly suited for comparison with observational data. Also important in our case is the fact that the post-Newtonian terms in the context of a parametrized post-Newtonian (PPN) expansion Will_book ; Will_exp ; Turyshev are expressed in these coordinates. Our choice therefore simplifies the job of comparing our results including a finite photon mass to the PPN expressions containing Eddinton’s $\gamma$ parameter, cf. App. A.
In Sec. II.1, as a warm up, we derive the deflection angle for a massive photon passing close to the Sun. Next, in Sec. II.2, we determine the gravitational time delay. In order to capture both gravitational and curvature effects, we will use the actual geodesic equations, obtained in App. B for convenience.
II.1 Deflection angle
One of the three classic tests of Einstein’s general relativity is the deflection of light by a massive object, which turns out to be twice as large as that derived from Newtonian mechanics: $\delta_{\rm E}=2\delta_{\rm N}=4GM_{\odot}/c^{2}r_{0}$, where $r_{0}$ is the distance of closest approach (for the Sun, it is the perihelion). Light, typically taken to be exactly massless, is just an extreme test case of the theory, since massive particles are obviously also affected by curvature. We are interested in the latter case, where photons are exceedingly light, but massive nevertheless.
In App. B we derived the geodesic equations and showed that we are allowed to constrain the trajectory to the equatorial plane ($\theta=\pi/2$) without loss of generality. Our target is to obtain an expression for $d\phi/dr$ as a function of the radial coordinate and, to this end, we start with eq. (42d) writing it as
$\dot{\phi}=\frac{K}{r^{2}\left(1+\frac{\mathcal{M}}{2r}\right)^{4}}\,,$ (3)
where $\mathcal{M}=GM_{\odot}/c^{2}\approx 1.5$ km.
The constant of integration $K$ may be determined by evaluating the constraint (42c) at the perihelion, $r=r_{0}$, where $\dot{r}=0$. By doing so, we find $K^{2}=c^{2}r_{0}^{2}F(\mathcal{M}/2r_{0})$ with $F(y)$ defined as
$F(y)=\left(1+y\right)^{4}\left[\frac{1}{\alpha^{2}}\frac{\left(1+y\right)^{2}}{\left(1-y\right)^{2}}-1\right]\,$ (4)
and $\alpha=m_{\gamma}c^{2}/E$. Plugging eq. (3) into the constraint $g_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu}=c^{2}$, cf. eq. (42c), we find
$\dot{r}=c\frac{r_{0}}{r}\frac{\sqrt{F(\mathcal{M}/2r_{0})}}{\left(1+\frac{\mathcal{M}}{2r}\right)^{4}}\left[\left(\frac{r}{r_{0}}\right)^{2}\frac{F(\mathcal{M}/2r)}{F(\mathcal{M}/2r_{0})}-1\right]^{1/2}\,,$ (5)
which duly satisfies $\dot{r}=0$ at $r=r_{0}$. Finally, dividing eq. (3) by eq. (5), we find
$\frac{d\phi}{dr}=\frac{1}{r}\left[\left(\frac{r}{r_{0}}\right)^{2}\frac{F(\mathcal{M}/2r)}{F(\mathcal{M}/2r_{0})}-1\right]^{-1/2}\,.$ (6)
We are interested in determining the total change in angle as the photon passes by the Sun. Noting that the orbit is symmetric around the perihelion, we only need to integrate the radial coordinate from $r_{0}$ to infinity and multiply by two. The total angle covered is $\Delta\phi=\pi+\delta$, with $\delta$ describing the gravity-induced deviation of the path from the unperturbed straight line. Changing variables to $q=r_{0}/r$ and letting $x=\mathcal{M}/2r_{0}$, we get
[…]
Noting that $x\ll 1$, we may expand the integrand and integrate term by term with the result
$|\Delta\phi|=\pi+4\left(\frac{2-\alpha^{2}}{1-\alpha^{2}}\right)x+\mathcal{O}(x^{2})\,.$ (8)
Since $\alpha^{2}\ll 1$ by assumption, we finally obtain
$\delta=\frac{4GM_{\odot}}{c^{2}r_{0}}\left(1+\frac{\alpha^{2}}{2}\right)\,$ (9)
to first post-Newtonian order. Clearly, the standard result is recovered in the massless limit and the photon mass appears squared, as expected Nieto . It is also clear that a finite photon mass acts to increase the deflection angle, which now depends on the energy of the incoming photon, meaning that different frequencies are deflected at different angles, a dispersive effect absent from the usual massless case.
Equation (9) matches the results from Refs. Accioly1 ; Accioly2 calculated via scattering methods, as well as those from a simplified calculation using a straight-line approximation, cf. App. A. In Ref. Accioly2 the authors use 43-GHz data from the Very-Long Baseline Array (VLBA) Fomalont and derive $m_{\gamma}\lesssim 3.5\times 10^{-5}\,{\rm eV/c^{2}}$, whereas in Ref. Accioly1 interferometric techniques are also used, but with 2-GHz radio waves Lebach , to obtain [^2] In Ref. Accioly1 the authors quote $m_{\gamma}\lesssim 10^{-40}\,{\rm g}\approx 5.6\times 10^{-8}\,{\rm eV/c^{2}}$. However, the deviation of the measured angles relative to Einstein’s result is $\Delta/\delta_{\rm E}\approx 8\times 10^{-4}$ Lebach , so that using the first equation in their section 3 we get $m_{\gamma}\lesssim 5.9\times 10^{-40}\,{\rm g}\approx 3.3\times 10^{-7}\,{\rm eV/c^{2}}$, a factor six worse than their stated limit. $m_{\gamma}\lesssim 3.3\times 10^{-7}\,{\rm eV/c^{2}}$. In Sec. III we will extract bounds complementary to those.
II.2 Gravitational time delay
One of the most interesting theoretical consequences of general relativity is that light signals propagating near a massive object will take a longer time to transit between source and receiver than what would be expected from Newtonian theory alone. This gravitational time delay was first calculated by Shapiro in 1964 in the context of radio signals sent from Earth towards Mercury or Venus close to superior conjunction Shapiro1 ; Shapiro2 ; Shapiro3 . In this section we calculate the Shapiro effect assuming that the signal is composed of massive photons following a similar strategy as above in the study of the deflection angle.
Our objective is to determine the total (coordinate) time elapsed from emission of the signal on Earth until its arrival at the receiver. Therefore, we must find $dt/dr$ as a function of known parameters, in particular the heliocentric radial distances of the objects involved. Once again we constrain the movement to the equatorial plane. We do not need to go back to the geodesic equations: it suffices to divide the constraint equation (42c) by $\dot{r}$ to get
$\left(\frac{dt}{dr}\right)^{2}$ $=$ $\frac{\left(1+\frac{\mathcal{M}}{2r}\right)^{6}}{c^{2}\left(1-\frac{\mathcal{M}}{2r}\right)^{2}}\left[1+r^{2}\left(\frac{d\phi}{dr}\right)^{2}\right]$ (10)
$+$ $\frac{\left(1+\frac{\mathcal{M}}{2r}\right)^{2}}{\left(1-\frac{\mathcal{M}}{2r}\right)^{2}}\frac{1}{\dot{r}^{2}}\,.$
As in Sec. II.1, it is useful to change variables to $q=r_{0}/r$, where $r_{0}$ is the shortest distance to the Sun, and define $x=\mathcal{M}/2r_{0}$, so that $\mathcal{M}/2r=qx$. With these changes and plugging in eqs. (5) and (6) for $\dot{r}$ and $d\phi/dr$, respectively, eq. (10) becomes
$\frac{dt}{dq}$ $=$ $\left(\frac{r_{0}}{c}\right)\frac{\left(1+qx\right)^{3}}{q^{2}\left(1-qx\right)}\frac{1}{\sqrt{H(q,x)}}$ (11)
$\times$ $\left[1+H(q,x)+\frac{\left(1+qx\right)^{4}}{q^{2}F(x)}\right]^{1/2}$
with $F(x)$ given by eq. (4) and $H(q,x)$ defined as
$H(q,x)=\frac{1}{q^{2}}\frac{F(qx)}{F(x)}-1\,.$ (12)
We must now compute the integral from the emitter or receiver of the signal at $q_{i}=r_{0}/r_{i}$ ($i=e,r$) to the perihelion at $q_{p}=1$. Before integrating, it is convenient to first expand the integrand in powers of $\alpha^{2}\ll 1$ and in $x=\mathcal{M}/2r_{0}\ll 1$. Keeping only the leading-order contributions, we get
$\Delta t_{\rm i\rightarrow p}$ $=$ $\frac{1}{c}\sqrt{r_{i}^{2}-r_{0}^{2}}\left(1+\frac{\alpha^{2}}{2}\right)$ (13)
$+$ $\frac{2GM_{\odot}}{c^{3}}\sqrt{\frac{r_{i}-r_{0}}{r_{i}+r_{0}}}\left(1+\alpha^{2}\right)$
$+$ $\frac{2GM_{\odot}}{c^{3}}\log\left(\frac{r_{i}+\sqrt{r_{i}^{2}-r_{0}^{2}}}{r_{0}}\right)\,.$
The one-way trip is composed of the path from emitter to receiver passing by the perihelion: $\Delta t_{\rm e\rightarrow p}+\Delta t_{\rm r\rightarrow p}$. Due to symmetry, the total coordinate time of the round trip is twice that, therefore
$\Delta t$ $=$ $\frac{2}{c}\left[\sqrt{r_{e}^{2}-r_{0}^{2}}+\sqrt{r_{r}^{2}-r_{0}^{2}}\right]\left(1+\frac{\alpha^{2}}{2}\right)$
$+$ $\frac{4GM_{\odot}}{c^{3}}\left[\sqrt{\frac{r_{e}-r_{0}}{r_{e}+r_{0}}}+\sqrt{\frac{r_{r}-r_{0}}{r_{r}+r_{0}}}\right]\left(1+\alpha^{2}\right)$
$+$ $\frac{4GM_{\odot}}{c^{3}}\log\left[\frac{\left(r_{e}+\sqrt{r_{e}^{2}-r_{0}^{2}}\right)\left(r_{r}+\sqrt{r_{r}^{2}-r_{0}^{2}}\right)}{r_{0}^{2}}\right]\,.$
The result above is not, however, what a clock on Earth would measure, since eq. (II.2) refers to the coordinate time as measured by a far-away observer. On Earth we must account for the gravitational influence of the Sun, as well as Earth’s translational motion. From the line element, cf. eq. (2), with $ds^{2}=c^{2}d\tau^{2}$ evaluated at $r=r_{e}$ (with $\dot{r}=0$) we have
$\left(\frac{\Delta\tau}{\Delta t}\right)^{2}\approx\left(1-\frac{2GM_{\odot}}{c^{2}r_{e}}\right)-\left(1+\frac{2GM_{\odot}}{c^{2}r_{e}}\right)r_{e}^{2}\frac{\dot{\phi}^{2}}{c^{2}}\,.$ (15)
Here $GM_{\odot}/c^{2}r_{e}\ll 1$ was used to expand the coefficients of the metric. Using the Newtonian approximation to Kepler’s third law, $r_{e}\dot{\phi}^{2}\approx GM_{\odot}/r_{e}^{2}$, we find that the proper time interval measured by a stationary clock on Earth is related to eq. (II.2) via
$\Delta\tau\approx\left(1-\frac{3GM_{\odot}}{2c^{2}r_{e}}\right)\Delta t\,,$ (16)
which gives a constant correction factor deviating from unity by [^3] The mean orbital radius of the Earth is $r_{e}=149\times 10^{6}$ km. $\approx 10^{-8}$.
With eqs. (II.2) and (16) the proper time elapsed between emission and arrival of the signal is finally
$\Delta\tau$ $=$ $\frac{2}{c}\left[\sqrt{r_{e}^{2}-r_{0}^{2}}+\sqrt{r_{r}^{2}-r_{0}^{2}}\right]\left(1+\frac{\alpha^{2}}{2}\right)$
-- $\frac{3GM_{\odot}}{c^{3}r_{e}}\left[\sqrt{r_{e}^{2}-r_{0}^{2}}+\sqrt{r_{r}^{2}-r_{0}^{2}}\right]$
$+$ $\frac{4GM_{\odot}}{c^{3}}\left[\sqrt{\frac{r_{e}-r_{0}}{r_{e}+r_{0}}}+\sqrt{\frac{r_{r}-r_{0}}{r_{r}+r_{0}}}\right]\left(1+\alpha^{2}\right)$
$+$ $\frac{4GM_{\odot}}{c^{3}}\log\left[\frac{\left(r_{e}+\sqrt{r_{e}^{2}-r_{0}^{2}}\right)\left(r_{r}+\sqrt{r_{r}^{2}-r_{0}^{2}}\right)}{r_{0}^{2}}\right]\,,$
where we neglected terms of order $\mathcal{O}(\alpha^{2}GM_{\odot}/c^{2}r_{e})$, as well as post-post-Newtonian corrections. It is worth keeping in mind that the gravitational contributions are of order $GM_{\odot}/c^{3}\approx 5\,{\rm\mu s}$. Note that finite-mass corrections affect only the first and third terms, but crucially not the phenomenologically most relevant logarithmic term. This will have important consequences when we extract limits on the photon mass in Sec. III.
Before closing this section, let us make a few general comments about our main result, eq. (II.2). The first term represents the Newtonian travel time in the absence of the Sun. This is effectively an Euclidean distance which cannot be measured directly due to the actual curvature of spacetime. The mass dependence of this term is expected, since the time it takes for a particle with normalized velocity $\beta=v/c$ to cover a distance $\ell$ is $(\ell/c)\beta^{-1}$. But in special relativity $E/m_{\gamma}c^{2}=(1-\beta^{2})^{-1/2}=\alpha^{-1}$, which can be rearranged as $\beta^{-1}=(1-\alpha^{2})^{-1/2}$. Expanding this expression yields $\beta^{-1}\approx 1+\alpha^{2}/2$, thus matching the mass-dependent factor in the first line of eq. (II.2). The other terms are general relativistic in nature and represent an extra travel time due to gravity. The gravitational time delay, or Shapiro effect, is defined as the non-Newtonian contribution to the total round-trip time: $\Delta\tau_{\rm delay}=\Delta\tau-\Delta\tau(M_{\odot}=0)$.
It is worth noting that in his first work Shapiro1 , Shapiro used Schwarzschild coordinates to compute the (proper) time delay in the straight-line approximation, but in his final analysis in Ref. Shapiro3 he mentions the use of harmonic coordinates and takes only the logarithm term as representative of the gravitational time delay. This is sensible, since this term is the most sensitive to the perihelion distance, $r_{0}$, which is the only time-dependent variable. Moreover, the logarithm term is the same in the simplified straight-line approximation, cf. App. A.1.
III Experimental sensitivities
In Sec. II.2 we obtained the correction to the two-way travel time of a signal as measured by an observer on Earth, mediated by massive photons, to first post-Newtonian order. Now we would like to use this result, together with current and future limits on deviations from standard general relativity (where the photon is massless), to set upper bounds on a putative rest mass of the photon.
Generally, the Shapiro effect refers to the non-Newtonian contribution to the round-trip time difference. In the context of tests of general relativity the gravitational time delay is expressed in the parameterized post-Newtonian (PPN) metric in terms of Eddington’s $\gamma$ parameter Poisson ; Will_book ; Will_exp , which characterizes curvature effects. Incidentally, it is also relevant in analyses of the deflection angle, cf. Sec. II.1.
As mentioned at the end of Sec. II.2, the logarithmic term is characteristic of the Shapiro effect due to its stronger dependence on the perihelion distance. Accordingly, the PPN expression for the time delay is usually quoted in the form of eq. (31), which was derived using a straight-line approximation for the light path (see discussion in App. A). However, using the actual geodesic we get eq. (37) and applying the correction factor (16) we find
$\Delta\tau_{\rm delay}^{\rm PPN}$ $=$ $\frac{4GM_{\odot}}{c^{3}}\left(\frac{1+\gamma}{2}\right)\left[2+\log\left(\frac{4r_{e}r_{r}}{r_{0}^{2}}\right)\right]$ (18)
-- $\frac{3GM_{\odot}}{c^{3}}\left(1+\frac{r_{r}}{r_{e}}\right)\,,$
where we assumed that $r_{0}\ll r_{e},r_{r}$, which is valid in all cases analysed here. Clearly, the logarithmic term is not the only contribution at post-Newtonian order. This expression will be our standard for comparisons with measurements of Eddington’s $\gamma$ parameter.
III.1 Bound from the Mariner 6 and 7 missions
The ten Mariner spacecraft were part of a program to study the inner solar system. In particular, Mariner 6 and 7 completed their flyby of Mars in 1969, whereupon they analysed its atmosphere and surface, relaying data back to Earth using 2.2 GHz (S-band) radio signals. Soon thereafter, the spacecraft entered heliocentric orbits.
In what follows we assume that the signals from emitter to receiver and back have approximately the same frequency. Subtracting the (unmeasurable) Newtonian light time from eq. (II.2) and comparing with eq. (18) we are able to express the deviation from general relativity, parametrized by $|\gamma-1|$, in terms of the photon mass as
$|\gamma-1|=\frac{4\alpha^{2}}{2+\log\left(\frac{4r_{e}r_{r}}{r_{0}^{2}}\right)}\,.$ (19)
Shorter perihelion distances increase the denominator, thereby reducing $|\gamma-1|$. This is sensible, since in this case light rays experience stronger gravitational fields and deviations from general relativity may be probed to better accuracy. On the other hand, this also means that our bounds will be weaker at shorter perihelion distances, a direct consequence of the fact that the logarithmic term is not affected by the finite photon mass, cf. eq. (II.2).
At this point it is useful to write $\alpha=m_{\gamma}c^{2}/E$ more explicitly as
$\alpha=2.4\times 10^{-9}\,\left(\frac{m_{\gamma}}{10^{-14}\,{\rm eV/c^{2}}}\right)\left(\frac{1\,{\rm GHz}}{f}\right)\,,$ (20)
where we used the quantum-mechanical relation $E=hf$. Assuming that no clear deviation from general relativity is found, any signs of new physics must lie below the uncertainty in $\gamma$, which is encoded in $|\gamma-1|$. Thus, using eq. (20) we have the following sensitivity
$m_{\gamma}$ $\lesssim$ $2.1\times 10^{-9}\,{\rm eV/c^{2}}\left(\frac{|\gamma-1|}{10^{-6}}\right)^{1/2}\left(\frac{f}{1\,{\rm GHz}}\right)$ (21)
$\times$ $\sqrt{2+\log\left(\frac{4r_{e}r_{r}}{r_{0}^{2}}\right)}\,.$
The behavior of eq. (21) as a function of the perihelion distance $r_{0}$ is displayed in Fig. 1.
Figure 1: The sensitivity (21) for $r_{e}=r_{r}=1$ AU, $|\gamma-1|=10^{-6}$ and $f=1$ GHz as a function of the perihelion distance $r_{0}$ in units of the solar radius. The slope is typical of the Shapiro time delay, cf eq. (II.2).
According to Ref. Mariner1 , superior solar conjunction for Mariner 6 and 7 happened around mid 1970 when the Earth-spacecraft distances were $402\times 10^{6}$ km (Mariner 6) and $386\times 10^{6}$ km (Mariner 7) with Sun-Earth-spacecraft angles of $0.95^{\circ}$ and $1.56^{\circ}$, respectively. This implies [^4] We use the Euclidean law of cosines as a good approximation. that $r_{r}\approx 253\times 10^{6}$ km and $r_{0}\approx 3.6R_{\odot}$ for Mariner 6, and $r_{r}\approx 237\times 10^{6}$ km and $r_{0}\approx 5.9R_{\odot}$ for Mariner 7. With the quoted uncertainty of $3\%$ on $\gamma$, we may use eq. (21) to calculate the upper bounds on the photon mass based on the data from each spacecraft. As it turns out, the results are very similar, differing only by $\approx 5\%$, so we take the weakest (from Mariner 6) as our conservative estimate, namely
$m_{\gamma}\lesssim 2.8\times 10^{-6}\,{\rm eV/c^{2}}\,.$ (22)
III.2 Bound from the Viking 1 and 2 missions
The Viking 1 and 2 missions, launched in 1975, had the detailed study of the Martian surface and atmosphere, and the search for signs of life as their main scientific goals. During their operation (at $r_{r}=1.52$ AU), the spacecraft worked with radio signals at $2.2$ GHz (S-band) with impact parameters as small as $r_{0}\approx 1.5R_{\odot}$ close to superior conjunction in late 1976 Viking ; Viking2 . The attained precision of $\approx 10$ ns led to a determination of $|\gamma-1|$ to within $0.2\%$ Viking3 . With this, eq. (21) gives us
$m_{\gamma}\lesssim 7.7\times 10^{-7}\,{\rm eV/c^{2}}\,,$ (23)
slightly improving the bounds from Mariner 6 and 7.
III.3 Bounds from the Cassini-Huygens mission
The best measurement of the Shapiro effect to date was performed using the Cassini spacecraft on its way to Saturn. During its superior conjunction in mid 2002, the spacecraft was at a distance $r_{r}=8.43$ AU and the electromagnetic signal reached its closest approach to the Sun at $r_{0}=1.6\,R_{\odot}$. The radio link was setup so that essentially two carrier signals were sent to and from the spacecraft. The uplink had $f_{u}=7.2$ GHz in the X-band and $f_{u}=34$ GHz in the Ka-band, whereas the downlink had $f_{d}=8.4$ GHz in the X-band and $f_{d}=32$ GHz in the Ka-band. The possibility of using carriers with different frequencies (in particular the Ka-band) allowed the almost complete elimination of coronal effects. Doppler-tracking of the signals via the fractional frequency shifts couuld be used to constrain $\gamma$ to unprecedented accuracy: $|\gamma-1|=2.3\times 10^{-5}$ Cassini_1 ; Cassini_2 .
To simplify the analysis, we will consider the X- and Ka-bands separately. This is also meaningful, as the finite photon mass induces dispersive effects. Since the up- and downlink frequencies are not the same, we are not allowed to use eq. (21). Instead, we must return to eq. (13) and treat the in- and outbound paths differently, assigning to each its frequency. Doing that we find
$\Delta\tau_{\rm delay}$ $=$ $\frac{4GM_{\odot}}{c^{3}}\left[\left(2+\alpha_{\rm u}^{2}+\alpha_{\rm d}^{2}\right)+\log\left(\frac{4r_{e}r_{c}}{r_{0}^{2}}\right)\right]$ (24)
-- $\frac{3GM_{\odot}}{c^{3}}\left(1+\frac{r_{c}}{r_{e}}\right)\,,$
where $\alpha_{\rm u,d}$ indicate the parameter (20) evaluated at frequency $f_{\rm u,d}$. This expression may now be compared with eq. (18), giving us the same sensitivity as in eq. (21), but with the substitution
$f\rightarrow\frac{\sqrt{2}f_{u}f_{d}}{\sqrt{f_{u}^{2}+f_{d}^{2}}}$ (25)
with $f_{u,d}$ in GHz. Using the values quoted above with eqs. (25) and (21), we finally obtain
[…]
The bound (26a) represents a 2.6-fold improvement over the bound from the Viking 1 and 2 data, cf. eq. (23).
III.4 Future perspectives
In the previous section we explored measurements of the gravitational time delay performed by past missions; the last one, Cassini-Huygens, was in the years 2000. Fortunately, there are a few promising proposals, either already underway or planned, that aim at significantly improving the current limits on $\gamma$, among other less well constrained PPN parameters Turyshev . A few examples are the Mercury Orbiter Radioscience Experiment (MORE) MORE ; MORE_1 and the Gravitational Time Delay Mission (GTDM) GTDM which, if all technical requirements are properly met, will be able to measure $|\gamma-1|$ at much higher accuracies than attained by the Cassini experiment. Their key parameters, along with those of the missions already discussed, are summarized in Table 1.
The MORE, which is on board of the BepiColombo mission launched in 2018 and planned to start orbiting Mercury in 2025 ($r_{r}=0.39$ AU), was designed to provide accurate tracking of the Mercury Planetary Orbiter MORE_1 . The mission aims at studying Mercury’s gravitational field to better understand its internal structure, as well as to perform precision tests of general relativity. For the latter, a similar strategy as for the Cassini mission is followed and two carriers are used for the up- and downlinks: X-band (7.2/8.4 GHz) and Ka-band (32/34 GHz) MORE_1 . With much improved ranging accuracy, it is projected that $|\gamma-1|\lesssim 2\times 10^{-6}$ will be achieved. Plugging these values in eq. (25) we get (assuming $r_{0}=2R_{\odot}$)
[…]
As in the case of the Cassini data, the lower frequency delivers the best results.
$f\,\left[{\rm GHz}\right]$ $r_{e}\,\left[{\rm AU}\right]$ $r_{r}\,\left[{\rm AU}\right]$ $r_{0}\,\left[R_{\odot}\right]$ $|\gamma-1|$
Mariner 6 $2.2$ $1$ $1.7$ $3.6$ $3\times 10^{-2}$bigstrut[t]
Mariner 7 $2.2$ $1$ $1.6$ $5.9$ $3\times 10^{-2}$bigstrut[t]
Viking 1,2 $2.2$ $1$ $1.52$ $1.5$ $2\times 10^{-3}$bigstrut[t]
Cassini 7.2-8.4/32-34 $1$ $8.43$ 1.6 $2.3\times 10^{-5}$bigstrut[t]
MORE 7.2-8.4/32-34 $\approx 1$ $0.39$ $2$ $2\times 10^{-6}$bigstrut[t]
GTDM $2.8\times 10^{5}$ $\approx 1$ $\approx 1$ $2$ $2\times 10^{-8}$bigstrut[t]
Table 1: Key parameters from selected future space-based missions aiming at improving the accuracy in measurements of Eddington’s $\gamma$ parameter (via the Shapiro effect). The data on the Mariner, Viking and Cassini missions are displayed for completeness. For MORE and GTDM we assume $r_{0}=2R_{\odot}$ when the respective transponder spacecraft are closest to superior conjunction; furthermore, the respective emitter spacecraft will be at the L1 Lagrange point $\approx 1.5\times 10^{6}$ km away from Earth in the direction of the Sun, thus $r_{e}\approx 1$ AU.
The GTDM GTDM will use interferometric techniques with $1064$ nm (282 THz) lasers to measure time delays. In fact, lasers are the current golden standard for ranging due to improved accuray and relative compactness of emitter and receiver systems Dirkx . The GTDM will send signals between two drag-free spacecraft – one located near the L1 Lagrange point between Earth and the Sun and the other in orbit at a heliocentric distance of $\approx 1$ AU – to measure time delays to within $0.4$ ps Turyshev . The expected accuracy will be enough to determine $|\gamma-1|$ at the level of $2\times 10^{-8}$, which in turn could be used to place the upper bound (assuming $r_{0}=2R_{\odot}$)
$m_{\gamma}\lesssim 3.0\times 10^{-4}\,{\rm eV/c^{2}}\,.$ (28)
This is significantly worse than the limits extracted from the data of much older missions or the projected bounds from MORE, despite of the higher accuracy in determining $\gamma$ from the GTDM. This is a direct consequence of the use of a high-frequency optical laser, which suppresses the sensitivity by a relative factor of $\approx 10^{5}$.
Figure 2: Bounds on the photon mass versus reported accuracy in the determination of Eddington’s $\gamma$ parameter (expressed as its deviation from unity). In blue are the bounds extracted from data of already completed missions; in red are projected sensitivities based on the design parameters of future time-delay measurements (cf. Table 1).
IV Concluding remarks
In this paper we analysed the gravitational time delay when the electromagnetic signal is composed of very light, but nonetheless massive, photons. Using heliocentric isotropic coordinates we worked out the equations of motion and, following the actual geodesic path, we obtained the two-way (proper) time delay to first post-Newtonian order, cf. eq. (II.2). By comparing this result with the corresponding PPN expression containing Eddington’s $\gamma$ parameter, eq. (18), we were able to extract upper bounds on the photon mass using the actual or projected limits on $|\gamma-1|$ from past and future time-delay measurements. The limits are summarized in Fig. 2.
From the standard PPN expression (18), we see that the time delay due to gravity increases as the light ray gets closer to the Sun (i.e., as $r_{0}$ decreases). Nonetheless, the sensitivity (21) does not benefit from this, as is evident from Fig. 1, where we see the emblematic logarithmic slope of the Shapiro effect acting to worsen the bound for smaller perihelion distances. This somewhat counter intuitive behavior is due to the fact that finite-mass effects, encoded in $\alpha=m_{\gamma}c^{2}/E$, modify only the unobservable Newtonian term and the almost constant post-Newtonian contribution, but not the logarithmic term in eq. (II.2); the correction to go from coordinate to proper time does not matter here. Since the logarithmic term is responsible for the actual time dependence of the Shapiro effect, getting closer to the solar limb means that the $\alpha$-dependent correction is suppressed, thereby worsening the limits on the photon mass.
The upper bounds stemming from data of the Cassini mission represent our best result, reaching the level of $m_{\gamma}\lesssim 3\times 10^{-7}\,{\rm eV/c^{2}}$, cf. eq. (26a). If the technical requirements and mission objectives of the MORE are fulfilled the limit may improve by a factor of $\sim 4$, cf. eq. (27a). Quite generally, our limits improve with decreasing frequency, an effect also reported elsewhere Accioly1 ; Accioly2 ; LSW . This is expected, since smaller masses are related to larger de Broglie wavelengths, meaning that low-frequency phenomena are better probes to finite-mass effects.
As discussed at the end of Sec. II.1, bounds of similar magnitude were reported in Refs. Accioly1 ; Accioly2 by using measurements of the deflection angle of $\sim 2-43$ GHz radio waves. Though the frequencies used were roughly the same, the accuracy in the determination of $\gamma$ were lower ($\sim 10^{-3}-10^{-4}$ Fomalont ; Lebach ) than what we have used. Nevertheless, their best upper bound is at the same level as ours from the Cassini data, cf. eq. (26a). The issue is, once again, which terms of the standard massless expression are modified by the photon mass. In the case of the deflection angle we have a deviation from general relativity $\sim\delta_{\rm E}\alpha^{2}$, cf. eq. (9). Therefore, any experimental bound will be directly proportional to the photon mass (squared), improving its sensitivity with even modest accuracies in $\gamma$. The situation with the Shapiro effect is unfortunately different, since the main phenomenological contribution, the logarithmic term in eq. (II.2), is not affected by finite-mass corrections. Consequently, the derived limits are not as strong as those from less precise measurements of the bending of radio waves by the Sun (see also the discussion below eq. (19)).
Future space-based missions will likely employ laser signals as a standard, which makes sense since optical frequencies are less impacted than radio by plasma effects in the solar corona, as these scale with $1/f^{2}$ Cassini_1 . Unfortunately, the use of optical lasers is less than ideal in the search for better limits on the photon mass, as is clear from the much better bounds from Cassini and MORE in comparison to the much more precise GTDM, cf. Fig. 2. From eq. (21) we see that radio links with frequencies at the MHz range would be sufficient to reach sensitivities around $m_{\gamma}\lesssim 10^{-13}\,{\rm eV/c^{2}}$ if $|\gamma-1|$ is determined at $10^{-8}$ accuracy, thereby approaching the realm of current bounds based on Earth-scale phenomena PDG ; Kroll ; Malta . In general, the lower the frequency, the better, and if technological limitations are overcome, time-delay measurements at sub-MHz-frequencies in space could potentially surpass other bounds obtained at solar-system scales Ryutov .
Acknowledgements.
We thank J.A. Helayël and Gustavo P. de Brito for reading the manuscript. P.C.M. is indebted to Marina and Karoline Selbach for insightful discussions. This work was financed in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brasil (CAPES) - Finance Code 001. C.A.D.Z. is partially supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) under the grant no. 310703/2021-2, and by Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro (Faperj) under Grant no. E-26/201.447/2021 (Programa Jovem Cientista do Nosso Estado).
Appendix A The massless case: Eddington’s $\gamma$ parameter and the PPN metric
In the main text we determined the gravitational time delay in the case of a massive photon that follows the geodesic path from emitter to receiver and back. This approach is often replaced by a simplified calculation using a straight line to approximate the true geodesic. In this appendix we would like to highlight the difference between the two, in particular for the standard case of a massless photon.
Our focus in this paper is to obtain limits on the photon mass by comparing the predictions with a finite mass to observational data. For this, it is useful to parametrize deviations from general relativity and the parametrized post-Newtonian (PPN) metric does just that Poisson ; Will_book ; Will_exp . The most relevant modification is that characterized by Eddington’s $\gamma$ parameter, which appears in the metric coefficients, cf. eq. (39), as ($\mathcal{M}=GM_{\odot}/c^{2}$)
$A(r)\approx 1-\frac{2\mathcal{M}}{r}\quad{\rm and}\quad B(r)\approx 1+\gamma\frac{2\mathcal{M}}{r}\,.$ (29)
In general relativity $\gamma=1$ and deviations may signal effects of new physics.
Here we compare the calculations based on the straight-line approximation with the more rigorous derivation based on the geodesic. From now on we deal exclusively with massless photons and, for the sake of simplicity, we do not apply the correction factor to convert coordinate time into proper time, cf. eq. (16).
A.1 Straight-line approximation
In the literature, most of the discussions of the Shapiro effect is done in terms of the usual Schhwarzschild coordinates using a straight line as an approximation for the curved path light takes. Let us constrain the motion to the $x-y$ plane, where emitter and receiver are at $(-x_{e},r_{0})$ and $(x_{r},r_{0})$, respectively, with $r_{0}$ being the point of closest approach.
The photon moves in the $x$-direction and the line element is $ds^{2}=0=A(r)c^{2}dt^{2}-B(r)dx^{2}$, so that, to leading order in $\mathcal{M}/r$,
$dt\approx\frac{1}{c}\left[1+(1+\gamma)\frac{\mathcal{M}}{r}\right]$ (30)
with $r=\sqrt{x^{2}+r_{0}^{2}}$. Integrating from the emitter to the receiver and multiplying by two we get
$\Delta t_{\rm line}=\frac{2(r_{e}+r_{r})}{c}+\frac{4GM_{\odot}}{c^{3}}\frac{(1+\gamma)}{2}\log\left(\frac{4r_{e}r_{r}}{r_{0}^{2}}\right)$ (31)
in the case where $r_{0}\ll x_{e},x_{r}\approx r_{e},r_{r}$. Note that the logarithmic term is the only relativistic correction to the Newtonian travel time.
A.2 Geodesic solution
Let us now repeat the calculation leading to eq. (II.2), but for a massless photon. We could start with the Lagrangean (38), but it is simpler to go another way. Since we are dealing with a massless photon, we cannot use the proper time as affine parameter. We may, however, use the coordinate time to parametrize the movement and use it as the line element such that
$dt^{2}=\frac{B(r)}{c^{2}A(r)}\left(dr^{2}+r^{2}d\phi^{2}\right)$ (32)
with $A(r),B(r)$ given by eq. (29).
Restricting the motion to the equatorial plane $\theta=\pi/2$, we may define a new Lagrangean
$L=\frac{B(r)}{c^{2}A(r)}\left(\dot{r}^{2}+r^{2}\dot{\phi}^{2}\right)\,.$ (33)
Again, $\phi$ is a cyclic coordinate and its associated Euler-Lagrange equation reads
$\dot{\phi}=\frac{K^{\prime}\left(1-\frac{2GM}{c^{2}r}\right)}{r^{2}\left(1+2\gamma\frac{GM}{c^{2}r}\right)}\,.$ (34)
The integration constant is determined by plugging eq. (34) into the metric (32) divided by $dt^{2}$ and evaluated at the perihelion, $r=r_{0}$, where $\dot{r}=0$. We find
$K^{\prime}=cr_{0}\left(\frac{1+2\gamma\frac{GM}{c^{2}r_{0}}}{1-2\frac{GM}{c^{2}r_{0}}}\right)^{1/2}\,.$ (35)
Inserting the results above in the constraint equation and isolating $\dot{r}$ we get (cf. eq. (29))
$c\frac{dt}{dr}=\sqrt{\frac{B(r)}{A(r)}}\Bigg{\{}1-\left(\frac{r_{0}}{r}\right)^{2}\frac{A(r)}{A(r_{0})}\frac{B(r_{0})}{B(r)}\Bigg{\}}^{-1/2}\,,$ (36)
which may be integrated similarly to eq. (10) by changing variables to $q=r_{0}/r$ and expanding in $2\mathcal{M}/r_{0}\ll 1$. The final result for the round trip with $r_{0}\ll x_{e},x_{r}\approx r_{e},r_{r}$ is
$\Delta t_{\rm geo}$ $=$ $\frac{2(r_{e}+r_{r})}{c}$ (37)
$+$ $\frac{4GM_{\odot}}{c^{3}}\frac{(1+\gamma)}{2}\left[2+\log\left(\frac{4r_{e}r_{r}}{r_{0}^{2}}\right)\right]\,.$
This result matches the one found by Weinberg Weinberg .
We may now contrast the PPN expressions for a massless photon obtained in the straight-line approximation, cf. eq. (31), with that determined via the geodesic, cf. eq. (37). Clearly, the Newtonian travel time is the same in both, as it should; the post-Newtonian terms, however, are different. In particular, for $\gamma=1$ we have a difference of $8GM_{\odot}/c^{3}\approx 40\,{\rm\mu s}$ between eqs. (31) and (37) that is purely due to the paths followed. Being practically constant (independent of $r_{0}$), this difference is not measurable in observations with varying perihelion distances, as was the case of the Cassini experiment Cassini_1 ; Cassini_2 , cf. Sec. III.3. It does, however, contribute to the total coordinate time elapsed, so we will employ eq. (37) as the standard PPN expression for the Shapiro effect in Sec. III.4.
Appendix B Geodesic equations in isotropic coordinates
[…]
$\frac{L}{m_{\gamma}}=A(r)c^{2}\dot{t}^{2}-B(r)\left[\dot{r}^{2}+r^{2}\left(\dot{\theta}^{2}+\sin^{2}\theta\dot{\phi}^{2}\right)\right]\,,$ (38)
where, as usual, the dot denotes differentiation with respect to proper time. Here the functions $A(r)$ and $B(r)$ are given by (cf. eq. (2))
$A(r)=\left(\frac{1-\frac{GM}{2c^{2}r}}{1+\frac{GM}{2c^{2}r}}\right)^{2}\quad{\rm and}\quad B(r)=\left(1+\frac{GM}{2c^{2}r}\right)^{4}\,.$ (39)
The Euler-Lagrange equations for the variables $q=\{t,r,\theta,\phi\}$ are
$2A(r)c^{2}\dot{t}$ $=$ $\epsilon\,,$ (40a)
$2B(r)\ddot{r}+A^{\prime}(r)c^{2}\dot{t}^{2}$ $=$ $B^{\prime}(r)\left[\dot{r}^{2}+r^{2}\left(\dot{\theta}^{2}+\sin^{2}\theta\dot{\phi}^{2}\right)\right]$ (40b)
-- $2B(r)r\left(\dot{\theta}^{2}+\sin^{2}\theta\dot{\phi}^{2}\right)\,,$
$B(r)r^{2}\sin\theta\cos\theta\dot{\phi}^{2}$ $=$ $\frac{d}{d\tau}\left[B(r)r^{2}\dot{\theta}\right]\,,$ (40c)
$-2B(r)r^{2}\sin^{2}\theta\dot{\phi}$ $=$ $\Lambda\,.$ (40d)
The primes in eq. (40b) indicate differentiation with respect to the argument of the function. Here $\epsilon$ and $\Lambda$ are constants stemming from the fact that $\{t,\phi\}$ are cyclical variables. These equations are not independent because of the constraint $g_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu}=c^{2}$, which reads
$A(r)c^{2}\dot{t}^{2}-B(r)\left[\dot{r}^{2}+r^{2}\left(\dot{\theta}^{2}+\sin^{2}\theta\dot{\phi}^{2}\right)\right]=c^{2}\,.$ (41)
Let us start by looking at eq. (40c) and consider movement in the equatorial plane $\theta=\pi/2$ with $\dot{\theta}=0$. These choices clearly satisfy this equation and imply that the trajectory will remain on this plane for all time, so we shall limit ourselves to this case. With this choice, the equations above reduce to
$2A(r)c^{2}\dot{t}$ $=$ $\epsilon\,,$ (42a)
$2B(r)\ddot{r}+A^{\prime}(r)c^{2}\dot{t}^{2}$ $=$ $B^{\prime}(r)\left(\dot{r}^{2}+r^{2}\dot{\phi}^{2}\right)$ (42b)
-- $2B(r)r\dot{\phi}^{2}\,,$
$A(r)c^{2}\dot{t}^{2}$ $=$ $B(r)\left(\dot{r}^{2}+r^{2}\dot{\phi}^{2}\right)+c^{2}\,,$ (42c)
$-2B(r)r^{2}\dot{\phi}$ $=$ $\Lambda\,.$ (42d)
If we move away from the Sun we have $A(r)\rightarrow 1$, so that eq. (42a) becomes $2c^{2}\dot{t}=\epsilon$, but now we are in flat Minkowski spacetime, where $\dot{t}=\gamma$, the Lorentz boost factor (not to be confused with Eddington’s $\gamma$ parameter), thus $2\gamma c^{2}=\epsilon$. We are therefore led to state that $\epsilon=2E/m_{\gamma}$ with $E$ being the total energy of the massive photon. Moving on to eq. (42d), at large distances, we get $B(r)\rightarrow 1$, so $-2r^{2}\dot{\phi}=\Lambda$. We recognize that, if $\Lambda=-2\ell/m_{\gamma}$, we recover the Newtonian expression for the (orbital) angular momentum, $\ell$.
With the constraint from eq. (42c) we may ignore eq. (42b) and, differentiating eqs. (42a), (42c) and (42d) with respect to the proper time we finally obtain the geodesic equations for the coordinates $\{t,r,\phi\}$:
$\ddot{t}$ $=$ $-\frac{2GM}{c^{2}r^{2}}\frac{\dot{r}\dot{t}}{\left(1-\frac{GM}{2c^{2}r}\right)\left(1+\frac{GM}{2c^{2}r}\right)}\,,$ (43a)
$\ddot{\phi}$ $=$ $-\frac{2}{r}\frac{\left(1-\frac{GM}{2c^{2}r}\right)}{\left(1+\frac{GM}{2c^{2}r}\right)}\dot{r}\dot{\phi}\,,$ (43b)
$\ddot{r}$ $=$ $-\frac{GM}{c^{2}r^{2}}\frac{\left(1-\frac{GM}{2c^{2}r}\right)}{\left(1+\frac{GM}{2c^{2}r}\right)^{7}}\dot{t}^{2}+\frac{\left(1-\frac{GM}{2c^{2}r}\right)}{\left(1+\frac{GM}{2c^{2}r}\right)}r\dot{\phi}^{2}$ (43c)
$+$ $\frac{GM}{c^{2}r^{2}}\frac{1}{\left(1+\frac{GM}{2c^{2}r}\right)}\dot{r}^{2}\,.$
Note that the equations above are independent of the constants of motion $\epsilon$ and $\Lambda$.
Appendix C Deflection angle in the straight-line approximation
In Sec. II.1 we based our calculation on the actual geodesics followed by a massive, ultra-relativistic particle when passing near a massive body. In this appendix we repeat the calculations under the assumption that the geodesics may be approximated by straight lines. For simplicity, we take the trajectory to lie on the $x-y$ plane. In this case we have $M_{\odot}$ at $y=0$ and the trajectory is taken to be parallel to the $y$-axis with $dx^{\beta}\approx(c/v,0,1,0)dy$, where $v$ is the velocity of the massive particle with 4-momentum $p^{\alpha}=(E/c,0,p^{y},0)$ Nieves .
Here we follow the approach from Refs. Ohanian ; Nieves and assume that the massive photon is moving along the positive $y$-direction with impact parameter $b$ along the $x$-axis. To first order in $GM_{\odot}/bc^{2}$, the total change in the massive photon’s 4-momentum is
[…]
where $h_{\alpha\beta}$ is the deviation of the metric from flat Minkowski spacetime. Assuming a weak field, we have $h_{00}=h_{ii}\approx-2GM_{\odot}/c^{2}r$, cf. eq. (2), and the change in 3-momentum in the $x$-direction is
$\Delta p_{x}=\frac{2GM_{\odot}}{c^{2}b}\left(\frac{E}{v}+p^{y}\right)\,.$ (45)
Energy and momentum are conserved and we are evaluating the deflection in asymptotically flat regions, where the metric is Minkowskian and the dispersion relation $E^{2}=m_{\gamma}^{2}c^{4}+(cp^{y})^{2}$ holds. Furthermore, energy and velocity are related via $E=m_{\gamma}c^{2}/\sqrt{1-v^{2}/c^{2}}$. Since $m_{\gamma}c^{2}\ll E$, we have
$\frac{E}{v}+p^{y}\approx\frac{2E}{c}\left(1+\frac{\alpha^{2}}{2}\right)\,,$ (46)
where $\alpha=m_{\gamma}c^{2}/E$. Noting that the deflection angle is given by $\delta=|\Delta p_{x}/p^{y}|\approx|\Delta p_{x}/(E/c)|$, we finally obtain
$\delta\approx\frac{4GM_{\odot}}{c^{2}b}\left(1+\frac{\alpha^{2}}{2}\right)\,,$ (47)
which matches Refs. Accioly1 ; Accioly2 , as well as our more rigorous result (9) derived using the full geodesic equations, provided we have $b\approx r_{0}$.